What is the acceleration of Jupiter as it moves in its circular orbit around the Sun? m/s2
Answer #1

let,

m is the mass of the Jupiter

M is the mass of the Sun=2*10^30 kg

r is the distance between Sun and Jupiter=7.785*10^11 m

v is the velocity of the Jupiter

now, in a circular motion

Fg=Fcentripetal

G*M*m/r^2=m*v^2/r

=====>

v^2=G*M/r

and

acceleration a=v^2/r=(G*M/r)/r

=(G*M/r^2)

=(6.67*10^-11*2*10^30)/(7.785*10^11)^2

=2.2*10^-4 m/sec^2